OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y + 2

OpenStudy (anonymous):

plz i need help

OpenStudy (anonymous):

i did not put the answers up because i need someone to show me how

OpenStudy (anonymous):

you have x + y = 6 x = y + 2 labeling them as eq A and B we get x + y = 6 ----->A x = y + 2------>B putting value if x from eq B in A we get y+2+y = 6 now from her find y as you find y put that value of y in eq B to get x thats it

OpenStudy (anonymous):

so i would do for the next one x+y+a?

OpenStudy (amoodarya):

where is x+y+a ??